mathslover (mathslover):

A particle experiences a constant accelerating force after starting from rest. If it travels a distance x in the first two seconds and a distance y in next two seconds, then : a) y = 4x b) y = 3x c) y = 2x d) y = x

mathslover (mathslover):

ok so it is y = x according to "me" .. @mukushla

mathslover (mathslover):

@UnkleRhaukus and @experimentX

OpenStudy (ghazi):

y=x

mathslover (mathslover):

That is totally wrong... I think ... (:P sorry I did it wrong I think this now ) the body is having "accelerating " force... and hence y can't be equal to x

mathslover (mathslover):

y > x

OpenStudy (ghazi):

distance traveled in 2 seconds=x distance traveled in 1 sec = x/2 similarly for y and when you'll equate this you'll have x=y ...time interval is equal and speed is not given so if we consider speed to be constant then y=x

mathslover (mathslover):

but please answer my doubt that the body is having "accelerating" force and HENCE We can not say that x = y ...

OpenStudy (ghazi):

sorry didn't see that...so Y=2x

mathslover (mathslover):

\[\large{h = ut + \frac{1}{2} at^2}\] \[\large{x = \frac{1}{2}(a)(4)}\] \[\large{x = 2a}\]

OpenStudy (ghazi):

you forgot u

mathslover (mathslover):

@ghazi u = 0 (given)

mathslover (mathslover):

now similarly we have : \[\large{h = ut+ \frac{1}{2}at^2}\] \[\large{(x+y) = \frac{1}{2}(a)(16) }\] \[\large{x+y = 8a}\] \[\large{2a+y=8a}\] \[\large{y = 6a}\]

mathslover (mathslover):

x = 2a y = 6a x/y = 2a / 6a = 1/3 x = 3/y y = 3x

mathslover (mathslover):

is that completely right? or any mistake here

mathslover (mathslover):

Though must say that idea came in mind when u suggested x/2 @ghazi

OpenStudy (ghazi):

that seems fair :)

OpenStudy (ghazi):

but for the next 2 seconds wouldn't there be some initial velocity

mathslover (mathslover):

Note that I took x + y ... not y that's why u = 0

OpenStudy (ghazi):



mathslover (mathslover):